The sgortest distance between the line x-3/3=y/0=z/-4 and y axis is.
options :
  1. 1/5
  1. 1
  1. 0
  1. 12/5
  2.  

Grade:12

1 Answers

Khimraj
5 years ago
\vec v=(3+3t)\vec i+(0+0*t)\vec j+(0-4t)\vec k\\ y=(0,1,0)\\ y*v=(0,1,0)*(3,0,-4)=(-4,0,-3)\\ v-y=(3,-1,0)\\ |(v-y).(y*v)|/|y*v|=12/5
Hope it clears.